judy obtained 80 marks out of 120 in her end of term maths examination. what was her percentage mark.Show all working
Found 2 solutions by checkley79, stanbon:
Answer by checkley79(3341)

80/120=.67 or 67% grade.

Answer by stanbon(75887)

judy obtained 80 marks out of 120 in her end of term maths examination. what was her percentage mark.Show all working
-----
x/100 = 80/120
x = 100(2/3)
---
x = 66 2/3%
=======================
Cheers,
Stan H.